A company made the followijg estimates for a componnet they use:

annual usage 8,000

ordering cost. $ 20 per order

carrying cost. 10% of cost per unit

cost per unit $ 20

based on this estimates an EOQ OF 400, an expected total cost of $800 p.a were calculated. During the year the EOQ of 400 was used for reordering,but the actual usage of components tremed to be 20% higher at 9,600

required:

calculate actual total stock costs

Calculate what the total costs wouldnhsve been if the correct EOQ HAD BEEN USED?

find out how sensitive are totsl cost erros in the usage estimate?
Answer #1

TC at EOQ 400

holding cost +ordering cost = 400/2 x 2 +8000/400 x20 =800

Revised cost if 20% more units consumed = 400/2 x2 + 9600/400 x 20 =880

Correct EOQ as per new quantity

[ 2x9600x20/2]1/2 = 438

New cost with EOQ = 438

438/2 x2 + 9600/438 x20 = 876

Cost error rate per increase of 38 units in the order quantity = 880-876 =4

The cost error per 20% increase in quantity is 4 x100/880 = 0.45%